there62ee are 850 douglar fir and ponderosa pine trees in a secion of forest bought by sawn. the company paid an average of 300 for each douglas fir and 225 for each ponderosa pine. if the company paid 217,500 for the trees,how many of rach kind did the company buy.
Found 2 solutions by checkley71, grohrbach:
Answer by checkley71(8403)

THERE62EE????????????? NEED TO CORRECT THIS WORD(S)

Answer by grohrbach(2)

There are 850 Douglas fir and ponderosa pine trees in a section of forest bought by Sawz Logging Co. The company paid an average of $300 for each Douglas fir and $225 for each ponderosa pine. If the company paid $217,500 for the trees, how many of each kind did the company buy?
x = number of Douglas fir trees
850 – x = number of Ponderosa pine trees
300x + 225(850 - x) = 217,500
300x + 191,250 - 225x = 217,500
75x + 191,250 = 217,500
- 191,250 - 191,250
____________________________
75x = 26,250 divide by 75
___divide
75x
x = 350 Douglas fir trees
850 – x = 500 Ponderosa Pine trees